### v2

# Improving the (approximate) sequential probability ratio test by avoiding overshoot

Abstract
The sequential probability ratio test (SPRT) by Wald (1945) is a cornerstone of sequential analysis. Based on desired type-I, II error levels $\alpha,\beta\in(0,1)$, it stops when the likelihood ratio statistic crosses certain upper and lower thresholds, guaranteeing optimality of the expected sample size. However, these thresholds are not closed form and the test is often applied with approximate thresholds $(1-\beta)/\alpha$ and $\beta/(1-\alpha)$ (approximate SPRT). When $\beta>0$, this neither guarantees type I,II error control at $\alpha,\beta$ nor optimality. When $\beta=0$ (power-one SPRT), it guarantees type I error control at $\alpha$ that is in general conservative, and thus not optimal. The looseness in both cases is caused by overshoot: the test statistic overshoots the thresholds at the stopping time. One standard way to address this is to calculate the right thresholds numerically, but many papers and software packages do not do this. In this paper, we describe a different way to improve the approximate SPRT: we change the test statistic to avoid overshoot. Our technique uniformly improves power-one SPRTs $(\beta=0)$ for simple nulls and alternatives, or for one-sided nulls and alternatives in exponential families. When $\beta>0$, our techniques provide valid type I and type II error guarantees, while needing less samples than Wald’s approximated thresholds in all considered simulations. These improved sequential tests can also be used for deriving tighter parametric confidence sequences, and can be extended to nontrivial settings like sampling without replacement and conformal martingales.
1 Introduction
Suppose we observe a stream of i.i.d. data $X_{1},X_{2},\ldots$ and are interested in a sequential test for the simple testing problem
$H_{0}:X_{i}\sim\mathbb{P}_{0}\quad\text{vs.}\quad H_{1}:X_{i}\sim\mathbb{P}_{1}.$ (1)
The classical approach to this testing problem is the sequential probability ratio test (SPRT) introduced by Wald [36, 37]. To this regard, we define the likelihood ratio as
$\Lambda_{0}:=1\quad\text{and}\quad\Lambda_{t}:=\frac{\prod_{i=1}^{t}p_{1}(X_{i})}{\prod_{i=1}^{t}p_{0}(X_{i})}\ (t\in\mathbb{N}),$
where $p_{0}$ and $p_{1}$ are the densities with respect to some measure $\mu$ of $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$, respectively. At some time $t$, the SPRT rejects $H_{0}$ if
$\Lambda_{t}\geq\gamma_{1},$
accepts $H_{0}$ if
$\Lambda_{t}\leq\gamma_{0},$
and continues sampling if
$\gamma_{0}<\Lambda_{t}<\gamma_{1},$
where $0<\gamma_{0}<\gamma_{1}$. Let $\delta_{\gamma_{0},\gamma_{1}}$ be the test decision of the SPRT ($\delta_{\gamma_{0},\gamma_{1}}=1$, if $H_{0}$ is rejected and $\delta_{\gamma_{0},\gamma_{1}}=0$, if $H_{0}$ is accepted) and $\tau_{\gamma_{0},\gamma_{1}}$ the stopping time. Wald [38] showed that $\tau_{\gamma_{0},\gamma_{1}}$ is almost surely finite. Furthermore, Wald and Wolfowitz [39] proved that for any other sequential test $(\delta^{\prime},\tau^{\prime})$ it holds that
$\mathbb{P}_{0}(\delta^{\prime}=1)\leq\mathbb{P}_{0}(\delta_{\gamma_{0},\gamma_{1}}=1)\text{ and }\mathbb{P}_{1}(\delta^{\prime}=0)\leq\mathbb{P}_{1}(\delta_{\gamma_{0},\gamma_{1}}=0)\implies\mathbb{E}_{h}(\tau^{\prime})\geq\mathbb{E}_{h}(\tau_{\gamma_{0},\gamma_{1}})\ (h\in\{0,1\}).$
Hence, the SPRT minimizes the expected sample size among all other sequential test with the same or smaller type I and type II error. However, in practice it is difficult or even impossible to calculate $\mathbb{P}_{0}(\delta_{\gamma_{0},\gamma_{1}}=1)$ and $\mathbb{P}_{1}(\delta_{\gamma_{0},\gamma_{1}}=0)$ exactly [36, 27]. Therefore, we cannot choose $\gamma_{0}$ and $\gamma_{1}$ such that the type I error is equal to a desired level $\alpha\in(0,1)$ and the type II error is equal to a desired level $\beta\in(0,1)$. For this reason, Wald [36, 37] introduced approximations for $\gamma_{1}$ and $\gamma_{0}$. Let $R_{1}^{t}=\{x=(x_{1},\ldots,x_{t}):\Lambda_{t}\geq\gamma_{1},\gamma_{0}<\Lambda_{i}<\gamma_{1}\ \forall i<t\}$ and $p_{h}^{t}(x)=\prod_{j=1}^{t}p_{h}(x_{j})$, $h\in\{0,1\}$. Wald’s approximation uses the fact that
$\mathbb{P}_{0}(\delta_{\gamma_{0},\gamma_{1}}=1)=\sum_{t\in\mathbb{N}}\int\limits_{R_{1}^{t}}p_{0}^{t}(x)\mu(dx)=\sum_{t\in\mathbb{N}}\int\limits_{R_{1}^{t}}\frac{p_{0}^{t}(x)}{p_{1}^{t}(x)}p_{1}^{t}(x)\mu(dx)\leq\gamma_{1}^{-1}\sum_{t\in\mathbb{N}}\int\limits_{R_{1}^{t}}p_{1}^{t}(x)\mu(dx)=\gamma_{1}^{-1}\mathbb{P}_{1}(\delta_{\gamma_{0},\gamma_{1}}=1)$
and therefore $\gamma_{1}\leq\frac{1-\mathbb{P}_{1}(\delta_{\gamma_{0},\gamma_{1}}=0)}{\mathbb{P}_{0}(\delta_{\gamma_{0},\gamma_{1}}=1)}$. In the same manner, one can show that $\gamma_{0}\geq\frac{\mathbb{P}_{1}(\delta_{\gamma_{0},\gamma_{1}}=0)}{1-\mathbb{P}_{0}(\delta_{\gamma_{0},\gamma_{1}}=1)}$. Wald’s approximation is based on replacing the inequalities with equalities, meaning to set
$\gamma_{1}=\frac{1-\beta}{\alpha}\text{ and }\gamma_{0}=\frac{\beta}{1-\alpha}.$
However, this only implies that $\frac{1-\beta}{\alpha}=\frac{1-\mathbb{P}_{1}(\delta_{\gamma_{0},\gamma_{1}}=0)}{\mathbb{P}_{0}(\delta_{\gamma_{0},\gamma_{1}}=1)}$ and $\frac{\beta}{1-\alpha}=\frac{\mathbb{P}_{1}(\delta_{\gamma_{0},\gamma_{1}}=0)}{1-\mathbb{P}_{0}(\delta_{\gamma_{0},\gamma_{1}}=1)}$, if $\Lambda_{\tau}\in\{\gamma_{0},\gamma_{1}\}$. The term overshoot is used to refer to how much $\Lambda_{\tau}$ exceeds the boundaries of $\gamma_{0}$ and $\gamma_{1}$, so the preceding condition $\Lambda_{\tau}\in\{\gamma_{0},\gamma_{1}\}$ assumes that there is no overshoot. However, this is not true in most testing scenarios, as the likelihood ratio is usually strictly larger than $\gamma_{1}$ or strictly smaller than $\gamma_{0}$ at the stopping time. In general, it can be shown [36] that if $\gamma_{1}=\frac{1-\beta}{\alpha}$ and $\gamma_{0}=\frac{\beta}{1-\alpha}$, then
$\mathbb{P}_{0}(\delta_{\gamma_{0},\gamma_{1}}=1)+\mathbb{P}_{1}(\delta_{\gamma_{0},\gamma_{1}}=0)\leq\alpha+\beta.$ (2)
Hence, not both the type I and type II error are inflated when using Wald’s approximation. Nevertheless, in many trials it is required that at least the type I error is controlled at the prespecified level $\alpha$. Furthermore, even if the sum of type I and type II errors are controlled, it is not ensured that the SPRT with Wald’s approximation minimize the expected sample size, meaning more efficient procedures are possible. To summarize, these thresholds $\gamma_{1}=\frac{1-\beta}{\alpha}$ and $\gamma_{0}=\frac{\beta}{1-\alpha}$ are only approximate, and therefore neither guarantee type I error control nor optimal use of sample size, and yet they are used very frequently in practice (e.g. [28, 14, 4, 10]). In particular, note that the publicly available R packages “sprtt: Sequential Probability Ratio Tests Toolbox” [29] and “SPRT: Wald’s Sequential Probability Ratio Test” [2] for performing an SPRT apply these approximate boundaries (the R Package “SPRT” was archived and removed from CRAN in 2022).
The power-one SPRT.
Equation (2) shows that the SPRT with Wald’s approximation nearly controls the type I error if $\beta$ is small. Indeed, it provably controls the type I error for $\beta=0$, as we will recap in the following. In this case, the SPRT stops for rejection if $\Lambda_{t}\geq 1/\alpha$, but never stops for accepting $H_{0}$. It is still guaranteed that the SPRT stops almost surely after a finite number of samples if $H_{1}$ is true [38], and therefore is called power-one SPRT, in particular studied and espoused by Robbins [7].
More abstractly, a level-$\alpha$ sequential test of power one (in short, a power-one sequential test) is defined by a stopping time $\tau$ (at which you stop and reject the null) such that
$\mathbb{P}_{0}(\tau<\infty)\leq\alpha$ (3)
and
$\mathbb{P}_{1}(\tau<\infty)=1,$ (4)
where $\alpha\in(0,1)$ is a prespecified significance level.
To see that the power-one SPRT satisfies (3), note that $(\Lambda_{t})_{t\in\mathbb{N}_{0}}$ defines a test martingale for $H_{0}$ (in the following we denote a process just by $(\Lambda_{t})$). That means $(\Lambda_{t})$ is nonnegative, $\mathbb{E}_{0}[\Lambda_{0}]=1$ and $\mathbb{E}_{0}[\Lambda_{t}|\Lambda_{t-1}]=\Lambda_{t-1}$. We will also speak of test supermartingale, if the last equality is replaced by an inequality. Ville’s inequality [30] implies that
$\mathbb{P}_{0}(\exists t\in\mathbb{N}:\Lambda_{t}\geq b)\leq 1/b$ (5)
for all $b\geq 1$. Hence, the power-one SPRT $\tau_{b}:=\inf\{t\in\mathbb{N}:\Lambda_{t}\geq b\}$, $b\geq 1$, satisfies condition (3) for $b=1/\alpha$. Furthermore, the power-one SPRT uses the available samples optimally in the following sense [5]:
$\mathbb{P}_{0}(\tau^{\prime}<\infty)\leq\mathbb{P}_{0}(\tau_{b}<\infty)\implies\mathbb{E}_{1}(\tau^{\prime})\geq\mathbb{E}_{1}(\tau_{b}).$
Consequently, any other sequential test that controls the type I error at the same level the SPRT does, needs on average at least as many samples as the SPRT to make a rejection if the alternative is true.
However, as for the classical SPRT, it is computationally very inefficient, or even infeasible, to calculate $\mathbb{P}_{0}(\tau_{b}<\infty)$ exactly. Therefore, one usually relies on (5) to satisfy (3) and thus sets $b=1/\alpha$. But Ville’s inequality (5) only becomes an equality, if $(\Lambda_{t})$ is never strictly larger than $b$ and almost surely attains 00 or $b$ at some point [18]. This is not fulfilled for most SPRTs. Hence, the SPRT is not optimal for $b=1/\alpha$.
Outline of the paper and our contribution.
In this paper, we propose a general boosting approach to uniformly improve (super)martingale based sequential tests by avoiding an overshoot at $1/\alpha$ (Section 2). This means that our boosted sequential tests are never worse than the initial tests, but may stop earlier or are more likely to achieve a rejection. In Section 3, we use this to uniformly improve power-one SPRTs for simple and one-sided nulls and alternatives. We provide nontrivial extensions and applications of our method in Section 4. In Section 5, we demonstrate how our boosting approach should be applied if a stop for futility is to be included and type II error control is desired. The R code for reproducing the simulation results and calculating the boosting factors is provided in the GitHub repository https://github.com/fischer23/boosting_SPRT.
In Table 1 we compare Wald’s approximate power-one and two-sided SPRT with our boosted versions. In the power-one case, our SPRT provides the same guarantees as Wald’s approximated SPRT, but uniformly improves it in terms of sample size. For the two-sided case, our boosted SPRT provides valid type I and type II error control while Wald’s approximated SPRT does not. In addition, our two-sided SPRT outperformed Wald’s with respect to sample size in all considered simulation scenarios.
Sequential test Type I control Type II control Sample size
Wald’s approx. power-one SPRT power-one Not optimal
Our boosted power-one SPRT power-one Not optimal, but never stops later than Wald’s approx. SPRT
Wald’s approx. two-sided SPRT - - Not optimal
Our boosted two-sided SPRT Not optimal, but often stops earlier than Wald’s approx. SPRT
Table 1: Comparison of the guarantees provided by Wald’s approximate power-one and two-sided SPRT, to our boosted SPRTs. Our boosted SPRT uniformly improves Wald’s approximated SPRT in terms of sample size in the power-one case, and often needs less samples in the two-sided case. In addition, our two-sided SPRT yields valid type I and type II error guarantee, while Wald’s approximated SPRT yields neither.
2 Improving test martingales by avoiding overshoot
The power-one SPRT martingale $(\Lambda_{t})$ wastes sample size by putting probability mass to values that are strictly larger than $1/\alpha$. In this section, we uniformly improve such (super)martingale based sequential tests by avoiding this overshoot.
Suppose we have some nonnegative process $(M_{t})$ with $\mathbb{E}_{0}[M_{0}]=1$ that is adapted to a filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}_{0}}$ (in the following we just write $(\mathcal{F}_{t})$), meaning $(\mathcal{F}_{t})$ is a nondecreasing sequence of sigma algebras and each $M_{t}$ is measurable with respect to $\mathcal{F}_{t}$. For example, one can think of $\mathcal{F}_{t}=\sigma(X_{1},\ldots,X_{t})$ as being generated by the observations, however, in some cases it is useful to coarsen the filtration or add randomization, as we shall later see with conformal martingales.
We define $L_{t}=M_{t}/M_{t-1}$ with the convention $/0=0$ 0 as the individual multiplicative factors of $(M_{t})$, so that
$M_{t}=M_{t-1}\cdot L_{t}=\prod_{i=1}^{t}L_{i}.$
Then $(M_{t})$ is a test supermartingale with respect to $(\mathcal{F}_{t})$ iff
$\mathbb{E}_{0}[L_{t}|\mathcal{F}_{t-1}]\leq 1\quad\text{for all }t\in\mathbb{N}.$ (6)
Hence, every test supermartingale can be decomposed into its individual factors. Furthermore, if factors $(L_{t})_{t\in\mathbb{N}}$ adapted to $(\mathcal{F}_{t})$ satisfying (6) are given, we can construct a new test supermartingale by $M_{t}=\prod_{i=1}^{t}L_{i}$.
Defining the stopping time
$\tau_{M}:=\inf\{t\in\mathbb{N}_{0}:M_{t}\geq 1/\alpha\}$
at which we reject the null, Ville’s inequality implies type-I error control: $\mathbb{P}_{0}(\tau_{M}<\infty)\leq\alpha$. Another useful property of test supermartingales is that
$\mathbb{E}_{0}[M_{\tau}]\leq 1\quad\text{for all stopping times }\tau,$ (7)
which follows from the optional stopping theorem. Mathematically, a stopping time $\tau$ is defined as an integer-valued random variable such that $\{\tau\leq t\}$ is measurable with respect to $\mathcal{F}_{t}$ for all $t\in\mathbb{N}$. A common interpretation is that a nonnegative random variable with expected value less or equal than one under the null hypothesis $H_{0}$ provides a fair bet against $H_{0}$ [24]. If “wealth” is accumulated by betting against $H_{0}$, there seems to be something odd with it. Thus, Property (7) implies that even if we stop at some time $\tau$ before rejecting the null hypothesis, a large value of $M_{\tau}$ can be seen as evidence against it.
Our approach is to increase the individual factors $L_{t}$ of a given test supermartingale $(M_{t})$ without violating its guarantee by avoiding an overshoot at $1/\alpha$. As a first step, we define the truncation function
$T_{\alpha}(x;M)=\begin{cases}x,&\text{ if }Mx\leq\frac{1}{\alpha}\\ \frac{1}{M\alpha},&\text{ if }Mx>\frac{1}{\alpha}.\end{cases}$ (8)
The truncation function ensures that $M\cdot T_{\alpha}(x;M)\leq 1/\alpha$ for all $x\in[0,\infty]$. Now define the process
$M_{t}^{\mathrm{trunc}}=M_{t-1}^{\mathrm{trunc}}\cdot T_{\alpha}(L_{i};M_{t-1}^{\mathrm{trunc}})=\prod_{i=1}^{t}T_{\alpha}(L_{i};M_{i-1}^{\mathrm{trunc}}).$ (9)
Note that, by construction, $M_{t}^{\mathrm{trunc}}$ never overshoots at $1/\alpha$. This can be verified by induction using the preceding property. Now, define
$\tau_{\mathrm{trunc}}:=\inf\{t\in\mathbb{N}_{0}:M_{t}^{\mathrm{trunc}}\geq 1/\alpha\}$
We now claim that
$\tau_{\mathrm{trunc}}=\tau_{M},$
meaning the sequential test based on $(M_{t}^{\mathrm{trunc}})$ makes the same decision and at the same time as the test based on $(M_{t})$. This is because for $t<\tau_{M}$, $M_{t}^{\mathrm{trunc}}=M_{t}$, meaning that the truncation has no effect if the threshold is not reached. At the very last step, when $M_{t}$ reaches (but may potentially exceed) $1/\alpha$, $M_{t}^{\mathrm{trunc}}$ simply equals $1/\alpha$, proving our claim.
However, since $T_{\alpha}(L_{t};M^{\mathrm{trunc}}_{t-1})\leq L_{t}$, we can potentially improve the factors $T_{\alpha}(L_{t};M^{\mathrm{trunc}}_{t-1})$ and thus the sequential test $\tau_{\mathrm{trunc}}$ while retaining valid error control. We propose to improve $L_{t}$ by multiplying it with a “boosting factor” $b_{t}\geq 1$ (the larger, the better). That means, at each step $t$, we calculate a boosting factor as large as possible such that
$\mathbb{E}_{0}[T_{\alpha}(b_{t}L_{t};M_{t-1}^{\mathrm{boost}})|\mathcal{F}_{t-1}]\leq 1,$ (10)
where $M_{t-1}^{\mathrm{boost}}=\prod_{i=1}^{t-1}T_{\alpha}(b_{i}L_{i};M_{i-1}^{\mathrm{boost}})$ and $(M_{t}^{\mathrm{boost}})$ is our boosted process that we use for testing. This procedure is summarized in Algorithm 1.
Algorithm 1 Improving test supermartingales by avoiding overshoot
Input: Test supermartingale $(M_{t})$ and individual significance level $\alpha$.
      Output: Boosted test supermartingale $(M_{t}^{\mathrm{boost}})$.
$M_{0}^{\mathrm{boost}}=1$
2:for $t=1,2,\ldots$ do
3:     $L_{t}=M_{t}/M_{t-1}$
4:     Choose $b_{t}\geq 1$ as large as possible such that $\mathbb{E}_{0}[T_{\alpha}(b_{t}L_{t};M_{t-1}^{\mathrm{boost}})|\mathcal{F}_{t-1}]\leq 1$
5:     Observe $L_{t}$
6:     $M_{t}^{\mathrm{boost}}=M_{t-1}^{\mathrm{boost}}T_{\alpha}(b_{t}L_{t};M_{t-1}^{\mathrm{boost}})$
7:end for
8:return $(M_{t}^{\mathrm{boost}})$
Theorem 2.1.
Let $(M_{t})$ be any test supermartingale and $(M_{t}^{\mathrm{boost}})$ be the boosted process obtained by Algorithm 1, yielding the sequential tests $\tau_{M}=\inf\{t\in\mathbb{N}:M_{t}\geq 1/\alpha\}$ and $\tau_{\mathrm{boost}}=\inf\{t\in\mathbb{N}:M_{t}^{\mathrm{boost}}\geq 1/\alpha\}$, respectively. Then $M_{t}^{\mathrm{boost}}\geq M_{t}$ for all $t<\tau_{M}$ and $\tau_{\mathrm{boost}}\leq\tau_{M}$ almost surely. Furthermore, $(M_{t}^{\mathrm{boost}})$ is a test supermartingale. Therefore, $\mathbb{P}_{0}(\tau_{\mathrm{boost}}<\infty)\leq\alpha$ and $\mathbb{E}_{0}[M_{\tau}^{\mathrm{boost}}]\leq 1$ for all stopping times $\tau$.
Proof.
We already noted that $\tau_{\mathrm{trunc}}=\tau_{M}$ and $M_{t}^{\mathrm{trunc}}=M_{t}$ for all $t<\tau_{M}$. Since $M_{t}^{\mathrm{boost}}\geq M_{t}^{\mathrm{trunc}}$ for all $t\in\mathbb{N}$, the first claim follows. Furthermore, since $\mathbb{E}_{0}[T_{\alpha}(b_{t}L_{t};M_{t-1}^{\mathrm{boost}})|\mathcal{F}_{t-1}]\leq 1$ for all $t\in\mathbb{N}$, it immediately follows that $(M_{t}^{\mathrm{boost}})$ is a test supermartingale.
Theorem 2.1 shows that Algorithm 1 provides a universal improvement technique for test supermartingales if a significance level $\alpha$ prespecified. Regardless of which test supermartingale $(M_{t})$ is put into Algorithm 1, the test based on the output process $(M_{t}^{\mathrm{boost}})$ always stops sooner ($\tau_{\mathrm{boost}}\leq\tau_{M}$). Furthermore, if we stop before the hypothesis is rejected, the boosted test martingale provides more evidence against the null hypothesis than the initial process ($M_{t}^{\mathrm{boost}}\geq M_{t}$ for all $t<\tau_{M}$). In the next section, we use Algorithm 1 to uniformly improve power-one SPRTs for simple and one-sided null and alternative hypotheses.
Remark 1.
The individual factors $L_{t}$ can be interpreted as sequential e-values recently explored in the literature [20, 33, 8]. A similar (but non-sequential) truncation approach like (8) was proposed by Wang and Ramdas [40] in the context of multiple testing. They also mentioned that in general we can use any “boosting function” $\psi_{t}:[0,\infty]\to[0,\infty]$ such that
$\mathbb{E}_{0}[T_{t}(\psi_{t}(L_{t});M_{t-1}^{\mathrm{boost}})|\mathcal{F}_{t-1}]\leq 1\quad\text{for all }t\in\mathbb{N}.$ (11)
However, recently Koning [16] proved that $\psi_{t}(x)=b_{t}x$, where $b_{t}$ is chosen such that (10) is satisfied with an equality, is the log-optimal choice. Therefore, we will stick to the boosting factor for the rest of the paper.
Remark 2.
Note that the boosted process $(M_{t}^{\mathrm{boost}})$ is in fact a test martingale if (10) holds with equality. Therefore, $(M_{t}^{\mathrm{boost}})$ can be interpreted as a likelihood ratio between $\mathbb{Q}$ and $\mathbb{P}_{0}$, where $\mathbb{Q}$ is adaptively determined ($\mathbb{Q}$ is a sub-probability distribution if $(M_{t}^{\mathrm{boost}})$ is only a test supermartingale).
3 Improving power-one SPRTs
In this section, we demonstrate how power-one SPRTs can be improved by the boosting technique proposed in Algorithm 1. SPRTs are particularly well-suited for our method, since each individual factor $L_{t}$ has the simple form $L_{t}=\lambda(X_{t})=\frac{p_{1}(X_{t})}{p_{0}(X_{t})}$. In addition, the distribution of $\lambda(X_{t})$ under the null hypothesis is usually known.
Let $\theta_{0},\theta_{1}\in\Theta$ with $\theta_{1}\geq\theta_{0}$ be two parameters, where $\Theta\subseteq\mathbb{R}$ is the parameter space of interest. Throughout this section, we assume that the likelihood ratio
$\lambda(x)=\frac{p_{\theta_{1}}(x)}{p_{\theta_{0}}(x)}\text{ is nondecreasing in }x,$ (12)
where $p_{\theta_{1}}$ and $p_{\theta_{0}}$ are densities with respect to some measure $\mu$. While this assumption is not required to apply our procedure, it helps to find general equations that determine the boosting factors and to test composite hypotheses.
An important class satisfying property (12) is the one-parameter exponential family, where each density is given by
$p_{\theta}(x)=h(x)\exp(T(x)\eta(\theta)-A(\theta))$
with $h$, $T$, $\eta$ and $A$ being known functions. It is easy to see that (12) is fulfilled for the sufficient statistic $T(x)$, if $\eta(\theta)$ is nondecreasing in $\theta$.
3.1 Simple null vs. simple alternative
In this subsection, we consider the simple testing problem
$H_{0}:\theta^{*}=\theta_{0}\quad\text{vs.}\quad H_{1}:\theta^{*}=\theta_{1},\ \theta_{1}>\theta_{0},$
where $\theta^{*}$ is the unknown parameter of interest. In order to derive general equations that determine the boosting factors, we assume that the likelihood ratio $\lambda:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$ has a generalized inverse $\lambda^{-1}:\mathbb{R}_{\geq 0}\rightarrow\mathcal{X}$ with the property $\lambda^{-1}(y)\geq x$ iff $\lambda(x)\leq y$ for all $x\in\mathcal{X}$ and $y\in\mathbb{R}_{\geq 0}$ (for example, this is fulfilled for $\lambda^{-1}(y)=\sup\{x\in\mathcal{X}:\lambda(x)\leq y\}$, if $\lambda$ is left-continuous).
With this, we can calculate
$\mathbb{E}_{\theta_{0}}[T_{\alpha}(b_{t}\lambda(X_{t});\Lambda_{t-1}^{\mathrm{boost}})|\Lambda_{t-1}^{\mathrm{boost}}]$
$=\mathbb{E}_{\theta_{0}}\left[b_{t}\lambda(X_{t})\mathbbm{1}\left\{b_{t}\lambda(X_{t})\leq\frac{1}{\alpha\Lambda_{t-1}^{\mathrm{boost}}}\right\}+\frac{1}{\Lambda_{t-1}^{\mathrm{boost}}\alpha}\mathbbm{1}\left\{b_{t}\lambda(X_{t})>\frac{1}{\alpha\Lambda_{t-1}^{\mathrm{boost}}}\right\}\Bigg{|}\Lambda_{t-1}^{\mathrm{boost}}\right]$
$=b_{t}\int_{x\leq\lambda^{-1}\left([b_{t}\alpha\Lambda_{t-1}^{\mathrm{boost}}]^{-1}\right)}\frac{p_{\theta_{1}}(x)}{p_{\theta_{0}}(x)}p_{\theta_{0}}(x)\mu(dx)+\frac{1}{\Lambda_{t-1}^{\mathrm{boost}}\alpha}\mathbb{P}_{\theta_{0}}\left(X_{t}>\lambda^{-1}\left(\frac{1}{b_{t}\alpha\Lambda_{t-1}^{\mathrm{boost}}}\right)\right)$
$=b_{t}F_{\theta_{1}}\left(\lambda^{-1}\left([b_{t}\alpha\Lambda_{t-1}^{\mathrm{boost}}]^{-1}\right)\right)+(\Lambda_{t-1}^{\mathrm{boost}}\alpha)^{-1}\left[1-F_{\theta_{0}}\left(\lambda^{-1}\left(\left[b_{t}\alpha\Lambda_{t-1}^{\mathrm{boost}}\right]^{-1}\right)\right)\right],$ (13)
where $F_{\theta_{h}}$ is the CDF of $\mathbb{P}_{\theta_{h}}$, $h\in\{0,1\}$ and $\Lambda_{t-1}^{\mathrm{boost}}=\prod_{i=1}^{t-1}b_{i}\lambda(X_{i})$. We can then set the equation equal to one and solve for $b_{t}$ numerically.
For example, in case of a simple Gaussian testing problem $H_{0}:X_{i}\sim\mathcal{N}(\mu_{0},1)$ vs. $H_{1}:X_{i}\sim\mathcal{N}(\mu_{1},1)$, $\mu_{1}-\mu_{0}=\delta>0$, we obtain
$\lambda(x)=\exp(\delta(x-\mu_{0})-\delta^{2}/2)\quad\text{and}\quad\lambda^{-1}(y)=\frac{\delta^{2}/2+\log(y)}{\delta}+\mu_{0}.$
Hence, (13) becomes
$b_{t}\left[\Phi\left(\frac{\log\left([b_{t}\alpha\Lambda_{t-1}^{\mathrm{boost}}]^{-1}\right)}{\delta}-\frac{\delta}{2}\right)\right]+(\Lambda_{t-1}^{\mathrm{boost}}\alpha)^{-1}\left[1-\Phi\left(\frac{\log([b_{t}\alpha\Lambda_{t-1}^{\mathrm{boost}}]^{-1})}{\delta}+\frac{\delta}{2}\right)\right],$ (14)
where $\Phi$ is the CDF of a standard Gaussian.
For example, suppose we test the above hypothesis for $\delta=2$ at level $\alpha=0.05$ and we are at the beginning of the testing process, meaning $\Lambda_{t-1}^{\mathrm{boost}}=\Lambda_{0}^{\mathrm{boost}}=1$. Then we obtain a boosting factor of $b_{t}=1.28$. If the current martingale value increases, the boosting factor increases as well, such that we obtain $b_{t}=1.55$ for $\Lambda_{t-1}^{\mathrm{boost}}=2$, but $b_{t}=1.14$ for $\Lambda_{t-1}^{\mathrm{boost}}=0.5$. In a similar manner, $b_{t}$ is increasing in $\delta$.
See Table 2 for boosting factors on a grid of values for $\delta$ and $\Lambda_{t-1}^{\mathrm{boost}}$. If $\delta$ and $\Lambda_{t-1}^{\mathrm{boost}}$ are both small, the boosting factors are close to one. The reason is that in this case the probability under $H_{0}$ of overshooting at $1/\alpha$ in the next step is close to zero. But also note that for small $\delta$ the expected sample size is usually large and this table only shows the boosting factors for one step. The total gain of boosting at some time could be calculated as the cumulative product of the boosting factors up to that time. The table also shows that the SPRT is very conservative for large $\delta$. For example, if $\delta=3$ and $\lambda(X_{1})=\lambda(X_{2})=1$, then $\Lambda_{2}=1$, but we would already have $\Lambda_{2}^{\mathrm{boost}}>20$ such that the boosted test could already reject the hypothesis.
$\delta$ $\Lambda_{t-1}^{\mathrm{boost}}$ $0.5$ $1$ $2$ $4$ $10$
$0.1$ $1$ $1$ $1$ $1$ $1.00001$
$0.5$ $1$ $1$ $1$ $1.00019$ $1.03019$
$1$ $1.00015$ $1.00157$ $1.01077$ $1.05386$ $1.37349$
$2$ $1.13931$ $1.27600$ $1.55046$ $2.17468$ $5.73972$
$3$ $2.45490$ $3.49439$ $5.72975$ $11.8255$ $68.1985$
Table 2: Boosting factors obtained for $\alpha=0.05$ in a simple Gaussian testing problem with mean difference $\delta$ and variance $1$.
[…]
$\mathbb{P}_{\theta_{0}}(\tau<\infty)=\mathbb{E}_{\theta_{1}}\left[\mathbbm{1}\{\tau<\infty\}\Lambda_{\tau}^{-1}\right],$ (15)
where $\Lambda_{\tau}^{-1}$ is the likelihood ratio of $\theta_{0}$ to $\theta_{1}$ at time $\tau$. With this, we can estimate the type I error by the inverse of the likelihood ratio at the stopping time while sampling under the alternative. Since $\tau$ is almost surely finite and usually much smaller under the alternative than under the null hypothesis, this reduces the required number of samples to obtain a good estimate of the type I error [26, 27].
The boosting approach saves on average $4\%$ to $12\%$ of the sample size compared to the SPRT in the considered scenarios, while the percentage increases with the strength of the signal. Furthermore, the boosted SPRT exhausts the significance level, while the SPRT is becoming increasingly conservative for stronger signals. In Appendix A we repeat this experiment for $\alpha=0.01$ instead of $\alpha=0.05$, which yields a similar result.
Figure 1: Comparison of sample size and type I error between the SPRT and the boosted SPRT in a simple Gaussian testing setup for different signal strengths and $\alpha=0.05$. The boosted SPRT saves $4-12\%$ of sample size compared to the SPRT and exhausts the significance level. The SPRT is becoming increasingly conservative for stronger signals.
3.2 Handling composite alternatives with the predictable plugin method
In practice, a concrete alternative is often difficult to specify, leading to testing problems of the form
$H_{0}:\theta^{*}=\theta_{0}\quad\text{vs.}\quad H_{1}:\theta^{*}>\theta_{0}.$
There are two approaches that are usually used to handle such composite alternatives in sequential testing. The method of mixtures [21] and the predictable plugin [36, 22, 23]. In the following, we will recap the latter, since it is better suited for our boosting technique.
Let $(\theta_{t})_{t\in\mathbb{N}}$ be a predictable process, meaning $\theta_{t}$ is measurable with respect to $\mathcal{F}_{t-1}$, such that $\theta_{t}\geq\theta_{0}$ for all $t\in\mathbb{N}$ and define
$\Lambda_{t}^{\text{plugin}}=\prod_{i=1}^{t}\frac{p_{\theta_{i}}(X_{i})}{p_{\theta_{0}}(X_{i})}.$ (16)
The idea of the predictable plugin method is to replace in the numerator of the likelihood ratio the density of a fixed alternative by an estimated density $p_{\theta_{t}}$ that is based on the previous data $X_{1},\ldots,X_{t-1}$. Since
$\mathbb{E}_{\theta_{0}}\left[\frac{p_{\theta_{i}}(X_{i})}{p_{\theta_{0}}(X_{i})}\bigg{|}\Lambda_{i-1}^{\text{plugin}}\right]=\int_{\mathcal{X}}\frac{p_{\theta_{i}}(x)}{p_{\theta_{0}}(x)}p_{\theta_{0}}(x)\mu(dx)=\int_{\mathcal{X}}p_{\theta_{i}}(x)\mu(dx)=1,$
it follows immediately that $(\Lambda_{t}^{\text{plugin}})$ defines a test martingale for $H_{0}$. Hence, the application of Algorithm 1 is straightforward. At each step $t\in\mathbb{N}$, we can calculate a boosting factor using (13) by just replacing $\theta_{1}$ with $\theta_{t}$.
A simple estimator used for the predictable plugin method is the smoothed maximum likelihood estimator (MLE)
$\theta_{t}=\max\left(\frac{\theta_{0}+\sum_{i=1}^{t-1}X_{i}}{t},\theta_{0}\right).$ (17)
In Figure 2 we repeated the experiment from Figure 1 in Section 3.1 but assumed that the alternative is not specified in advance using the predictable parameter in (17). Unsurprisingly, the sample size required by both methods is larger than when the alternative was prespecified. However, the gain in sample size achieved by our boosting method is even larger in the case of the composite alternative, as we were able to save $6\%$ to $14\%$ of the sample size compared to the usual SPRT.
Figure 2: Comparison of sample size between the SPRT with a predictable plugin and the boosted SPRT in a Gaussian testing setup with composite alternative for different signal strengths. The boosted SPRT saves $6-14\%$ of sample size compared to the SPRT.
3.3 Handling composite nulls in one-sided testing problems
An even more common testing setup in practice is
$H_{0}:\theta^{*}\leq\theta_{0}\quad\text{vs.}\quad H_{1}:\theta^{*}>\theta_{0}\quad(\theta_{0}\in\Theta),$
meaning we have a composite null hypothesis and a composite alternative. Robbins and Siegmund [22] showed that the process $Z_{t}$ defined in Equation (16) also provides a valid sequential test for the composite null hypothesis above. In the following, we show that the same holds for our boosted process.
Due to the monotone likelihood ratio property (12), we know that
$\mathbb{P}_{\theta_{0}}\left(\frac{p_{\theta_{t}}(X_{t})}{p_{\theta_{0}}(X_{t})}\geq x\right)\geq\mathbb{P}_{\theta}\left(\frac{p_{\theta_{t}}(X_{t})}{p_{\theta_{0}}(X_{t})}\geq x\right)\quad(\theta\leq\theta_{0},x\geq 0),$ (18)
where $\theta_{t}>\theta_{0}$. Writing $\lambda_{t}(X_{t})=\frac{p_{\theta_{t}}(X_{t})}{p_{\theta_{0}}(X_{t})}$, Equation (18) implies that
$\mathbb{P}_{\theta_{0}}\left(T_{\alpha}(b_{t}\lambda_{t}(X_{t});M)\geq x\right)\geq\mathbb{P}_{\theta}\left(T_{\alpha}(b_{t}\lambda_{t}(X_{t});M)\geq x\right)\quad(\theta\leq\theta_{0},x\geq 0)$
for any $\alpha\in(0,1)$, $M>0$ and $b_{t}\geq 1$. Hence, if $\mathbb{E}_{\theta_{0}}[T_{\alpha}(b_{t}\lambda_{t}(X_{t});M)|\Lambda_{t-1}^{\mathrm{boost}}]\leq 1$, we also have $\mathbb{E}_{\theta}[T_{\alpha}(b_{t}\lambda_{t}(X_{t});M)|\Lambda_{t-1}^{\mathrm{boost}}]\leq 1$ for all $\theta<\theta_{0}$. This implies that we can just use the likelihood ratio of $\theta_{t}$ vs. $\theta_{0}$ and derive the boosting factors by Algorithm 1 using $\mathbb{E}_{\theta_{0}}$, but the boosted test supermartingale will also be valid for all $\theta\leq\theta_{0}$.
To summarize, for the composite testing problem above we can do exactly the same as described in Section 3.2 for the simple null hypothesis while providing valid error control.
4 Nontrivial extensions and applications
In this section, we demonstrate some nontrivial extensions and applications of our boosting method. First, we show how it can be used for tighten confidence sequences in Section 4.1. Afterwards, we exemplify its use in sampling without replacement situations (Section 4.2) and with conformal martingales (Section 4.3) which are employed for testing the assumption of i.i.d. data.
4.1 Confidence sequences
A $(1-\alpha)$-level confidence sequence (CS) for some unknown parameter $\theta^{*}\in\Theta$ is a sequence of sets $(C_{t})_{t\in\mathbb{N}}$ such that
$\mathbb{P}_{\theta^{*}}(\exists t\in\mathbb{N}:\theta^{*}\notin C_{t})\leq\alpha.$ (19)
Let $\tau_{\theta}$, $\theta\in\Theta$, be a sequential test for $H_{0}^{\theta}:\theta^{*}=\theta$ with property (3). Then
$C_{t}=\{\theta\in\Theta:\tau_{\theta}>t\}$
is a CS for $\theta^{*}$. Let $(M_{t}^{\theta})$ be the test supermartingale for $H_{0}^{\theta}$ that defines $\tau_{\theta}$, meaning $\tau_{\theta}=\inf\{t\in\mathbb{N}:M_{t}^{\theta}\geq 1/\alpha\}$. It follows immediately that $(C_{t})_{t\in\mathbb{N}}$ can be boosted, meaning each $C_{t}$ shrinks in size while maintaining (19), by boosting each of the martingales $(M_{t}^{\theta})$, $\theta\in\Theta$, by Algorithm 1. However, the computational effort of this approach is extremely high or even infeasible.
Due to computational convenience and better interpretability, in practice one usually focuses on CSs that form an interval
$C_{t}=(l_{t},u_{t})$
for some lower bound $l_{t}$ and upper bound $u_{t}>l_{t}$. We will focus on the construction of a lower bound in the following. The two-sided CSs can then be obtained as the intersection of an upper and a lower $(1-\alpha/2)$-level confidence sequence. Hence, we are looking for a sequence $(l_{t})_{t\in\mathbb{N}}$ such that
$\mathbb{P}_{\theta^{*}}(\exists t\in\mathbb{N}:\theta^{*}<l_{t})\leq\alpha.$
In order to ensure that the resulting CS has this form, it is required that the martingales $(M_{t}^{\theta})_{\theta\in\Theta}$ are nonincreasing in $\theta$ for every $t\in\mathbb{N}$. Now note that each boosting factor $b_{t}^{\theta}$, $\theta\in\Theta$ and $t\in\mathbb{N}$, defined by Algorithm 1, is a nondecreasing function of the (boosted) supermartingale at the previous step $M_{t-1}^{\theta,\mathrm{boost}}$. Therefore, the boosting factors $(b_{t}^{\theta})_{\theta\in\Theta}$ are also nonincreasing in $\theta$, provided that the distribution of $(M_{t}^{\theta})$ under $H_{0}^{\theta}$ is the same for all $\theta\in\Theta$. This implies that also the boosted processes $(M_{t}^{\theta,\mathrm{boost}})_{\theta\in\Theta}$ are nonincreasing in $\theta$ for each $t$ such that the resulting CS provides indeed a lower bound. With this, the boosted CS can be easily determined computationally by setting $l_{t}$ at each step $t\in\mathbb{N}$ to be the smallest $\theta\in\Theta$ such that $M_{t}^{\theta,\mathrm{boost}}<1/\alpha$. In the following, we illustrate the construction of a boosted CS with a Gaussian example.
Example 1.
Suppose we look for a confidence sequence for the mean $\mu^{*}$ of i.i.d. Gaussians $X_{1},X_{2},\ldots$ with known variance one. For each $\mu\in\mathbb{R}$, we need to determine a sequential test for the null hypothesis
$H_{0}^{\mu}:\mu^{*}=\mu.$
Let $\mu\in\mathbb{R}$ be fixed and define $Z_{t}=X_{t}-\mu$ for all $t\in\mathbb{N}$ such that each $Z_{t}$ is standard normally distributed under $H_{0}^{\mu}$. This implies that
$\mathbb{E}_{\mu}[\exp(\delta Z_{t}-\delta^{2}/2)|\mathcal{F}_{t-1}]=1,$
where $\delta>0$ is some constant. Hence, the process $(M_{t}^{\mu})$, where
$M_{t}^{\mu}=\prod_{i=1}^{t}\exp(\delta Z_{i}-\delta^{2}/2)$
provides a test martingale for $H_{0}^{\mu}$. Furthermore, note that $M_{t}^{\mu}$ is decreasing in $\mu$ for each $t$, such that the family of test martingales $((M_{t}^{\mu}))_{\mu\in\mathbb{R}}$ can be used to find a lower bound $l_{t}$ for $\mu^{*}$. Indeed, we can calculate
$M_{t}^{\mu}<1/\alpha\Leftrightarrow\prod_{i=1}^{t}\exp(\delta(X_{i}-\mu)-\delta^{2}/2)<1/\alpha\Leftrightarrow\frac{1}{t}\sum_{i=1}^{t}X_{i}-\frac{\log(1/\alpha)}{n\delta}-\delta/2<\mu.$
Consequently, $(l_{t})_{t\in\mathbb{N}}$, where $l_{t}=\frac{1}{t}\sum_{i=1}^{t}X_{i}-\frac{\log(1/\alpha)}{n\delta}-\delta/2$, is a valid lower bound sequence for $\mu^{*}$. By setting $Z_{t}=\mu-X_{t}$ one can derive an upper bound, yielding the $(1-\alpha)$ confidence sequence
$\frac{1}{t}\sum_{i=1}^{t}X_{i}\ \pm\ \left[\frac{\log(2/\alpha)}{t\delta}+\delta/2\right].$ (20)
This CS was already proposed by Robbins and colleagues [6, 21]. As described above, one can obtain a boosted confidence sequence numerically by boosting each $(M_{t}^{\mu})$, $\mu\in\mathbb{R}$, separately and choosing the smallest/largest $\mu\in\mathbb{R}$ such that $M_{t}^{\mu,\mathrm{boost}}<1/\alpha$ for the lower/upper bound. In Figure 3 we compare the $(1-\alpha)$ lower confidence bound proposed by Robbins with our boosted improvement for $\mu^{*}=2$ (the bounds were obtained by averaging over $100$ trials), which shows that boosting can tighten the CS (20) significantly. We chose $\delta=\sqrt{8\log(1/\alpha)/n}$ for $n=50$, as recommended for obtaining a tight CS at time $n$ [11].
Note that Robbins proved that the CS (20) is valid for all 1-sub-Gaussian distributions. However, our boosting approach only ensures valid error control for Gaussian distributions with variance one. The reason is that
[…]
for $b_{t}>1$ and therefore it is not possible to handle all 1-sub-Gaussian distributions simultaneously. It is interesting that it is possible to exploit the information of the observations being exactly Gaussian (instead of just 1-sub-Gaussian), since previous works on confidence sequences did not make use of it [6, 21, 11, 12].
Figure 3: Comparison of the lower confidence bound (20) by Howard et al. [11] with its boosted improvement for $\mu^{*}=2$. Boosting improves the CS significantly.
4.2 Sampling without replacement
Suppose we have some finite population $\{x_{1},\ldots,x_{N}\}$ and draw samples
$X_{t}|\{X_{1},\ldots,X_{t-1}\}\sim\text{Uniform}(\{x_{1},\ldots,x_{N}\}\setminus\{X_{1},\ldots,X_{t-1}\})$
sequentially and without replacement with the goal of testing the hypothesis
$H_{0}:\mu^{*}\leq\mu_{0}\quad\text{vs.}\quad H_{1}:\mu^{*}>\mu_{0},$
where $\mu^{*}=\frac{1}{N}\sum_{i=1}^{N}x_{i}$, because the effort of drawing all samples is too high. In this section, we focus on binary variables $X_{i}$. For example, this scenario is encountered when auditing the outcome of an election [43] or when checking whether the permutation p-value is below the predefined level $\alpha$ [42]. In the following, we show how the state-of-the-art method can be improved by our boosting technique.
Waudby-Smith et al. [43] proposed the RiLACS algorithm to test $H_{0}$, which is based on the test martingale
$M_{t}=\prod_{i=1}^{t}(1+\lambda_{i}(X_{i}-C_{i}(\mu_{0}))),$ (21)
where $C_{i}(\mu_{0}))=\frac{N\mu_{0}-\sum_{j=1}^{i-1}X_{j}}{N-i+1}$ is the conditional mean of $X_{i}|X_{1},\ldots,X_{i-1}$ if the mean of $\{x_{1},\ldots,x_{N}\}$ is $\mu_{0}$ and $\lambda_{i}\in\left[0,\frac{1}{C_{i}(\mu_{0})}\right]$ is a hyperparameter measurable with respect to $\mathcal{F}_{i-1}$. Hence, each individual factor of $(M_{t})$ is just given by $L_{t}=1+\lambda_{i}(X_{i}-C_{i}(\mu_{0}))$. Under the null hypothesis and conditional on $\mathcal{F}_{t-1}$, we have that $L_{t}=1+\lambda_{i}(1-C_{i}(\mu_{0}))$ with probability at most $C_{i}(\mu_{0})$ and $L_{t}=1-\lambda_{i}C_{i}(\mu_{0})$ with probability at least $1-C_{i}(\mu_{0})$. This implies that we can determine valid boosting factors $b_{t}$ by
$b_{t}=\begin{cases}\frac{1-C_{i}(\mu_{0})/(M_{t-1}^{\mathrm{boost}}\alpha)}{(1-C_{i}(\mu_{0}))((1-\lambda_{i}C_{i}(\mu_{0}))},&\text{ if }M_{t-1}^{\mathrm{boost}}(1+\lambda_{i}(1-C_{i}(\mu_{0})))\geq 1/\alpha\\ 1,&\text{ otherwise.}\end{cases}$
If the alternative $H_{1}:\mu^{*}=\mu_{1}$ is known in advance, Waudby-Smith et al. [43] proposed to choose $\lambda_{i}=\min(1/C_{i}(\mu_{0}),2(2\mu_{1}-1))$. For example, this is the case in Risk-limiting audits (RLA) used for auditing the results of an election. In Figure 4, we compare the required sample size of RiLACS and its boosted improvement for $\mu_{0}=0.5$ and $\mu_{1}=0.55$ for different total number of samples $N$ at level $\alpha=0.01$. The gain in sample size obtained by boosting is small but visible. This is not surprising, since the RiLACS test martingale (21) only overshoots slightly at $1/\alpha$ due to the binary nature of the data.
Figure 4: Comparison of the required sample size by RILACS [43] and its boosted improvement in a sampling without replacement setting for different total number of samples. Boosting leads to a visible gain in sample size. However, due to the binary nature of the data, overshoot is not that big of a problem and therefore the gain in sample size is smaller than in Gaussian examples.
4.3 Conformal martingales
A common assumption in statistics has been that the data is generated independently from the same distribution, also called iid. (independent and identically distributed). In this subsection, we consider testing this assumption sequentially, meaning to test the null hypothesis
$H_{0}:X_{1},X_{2},\ldots\text{ are iid.}$.
From a single data sequence it is impossible to distinguish whether the data is iid or exchangeable [19], where $(X_{t})_{t\in\mathbb{N}}$ is called exchangeable if $(X_{1},\ldots,X_{t})$ has the same joint distribution as $(X_{\sigma(1)},\ldots,X_{\sigma(t)})$ for every $t\in\mathbb{N}$ and permutation $\sigma$ of the first $t$ indices. Hence, $H_{0}$ can be formulated equivalently as
$H_{0}:X_{1},X_{2},\ldots\text{ are exchangeable.}$.
Note that this is a highly composite null hypothesis. However, Vovk’s conformal martingale approach [31, 34, 35] allows to reduce $H_{0}$ to a simple null hypothesis. This makes it not only possible to test $H_{0}$ sequentially using a test martingale, but also to apply our boosting technique from Algorithm 1. We first recap the method by Vovk and then demonstrate how we can apply our boosting approach on top of it.
Let $A$ be a nonconformity measure, which maps observations $(x_{1},\ldots,x_{t})$ of any length to a sequence of numbers $(z_{1},\ldots,z_{t})\in\mathbb{R}^{t}$ such that for any $t$ and any permutation $\sigma$ it holds that
$A(x_{1},\ldots,x_{t})=(z_{1},\ldots,z_{t})\implies A(x_{\sigma(1)},\ldots,x_{\sigma(t)})=(z_{\sigma(1)},\ldots,z_{\sigma(t)}).$
The number $z_{i}$ can be interpreted as the “strangeness” of $x_{i}$. With this, for each $t\in\mathbb{N}$, one can calculate a “p-value” by
$p_{t}=\frac{|\{i\leq t:z_{i}>z_{t}\}|+\theta_{t}|\{i\leq t:z_{i}=z_{t}\}|}{t},$
where $(z_{1},\ldots,z_{t})=A(x_{1},\ldots,x_{t})$ and $\theta_{t}$ is randomly and independently sampled from $U[0,1]$. Vovk showed that the p-values $p_{1},p_{2},\ldots$ are iid from $U[0,1]$ under the null hypothesis $H_{0}$.
Vovk proposed to collect evidence against $H_{0}$ by forming a conformal martingale $(M_{t})$ as
$M_{t}=\prod_{i=1}^{t}f_{i}(p_{i}),$
[…]
A concrete betting function proposed by Vovk is
$f_{\kappa}(u)=\kappa u^{\kappa-1},\quad\kappa\in(0,1).$
Since we know that the p-values $p_{1},p_{2},\ldots$ are iid uniformly distributed on $U[0,1]$ under $H_{0}$, we can easily determine the distribution of $f_{\kappa}(p_{i})$ conditional on the previous p-values for every $\kappa$ such that Algorithm 1 can be applied. For example, $\alpha=0.05$, $M_{t-1}^{\mathrm{boost}}=1$ and $\kappa=0.5$ yield a boosting factor of $b_{t}=1.013$. The boosting factor is decreasing in $\kappa$ such that we obtain a boosting factor of $b_{t}=2.199$ for $\kappa=0.1$.
5 Challenging Wald’s two-sided approximate SPRT
While the previous sections focused on power-one SPRTs ($\beta=0$) where we only stop to reject the null but not to accept it, we now consider the case of $\beta>0$, where we also wish to stop if it is unlikely that we would reject the null if we continued testing. We will show how we can use a similar boosting technique as in Section 2 to construct powerful anytime-valid tests that allow to “stop for futility” (stop and accept the null).
Our main comparison point will be Wald’s approximate two-sided SPRT, which uses thresholds of $(1-\beta)/\alpha$ and $\beta/(1-\alpha)$. Importantly, out of the three desiderata of (a) type-I error control at $\alpha$, (b) type-II error control at $\beta$, (c) optimal sample size, these approximate thresholds guarantee none of the three properties. In this section, we will first develop a method with valid type-I error control (as a warm-up), and then one with valid type-I and type-II error control. The latter method often outperforms Wald’s approximate SPRT in the sense that while guaranteeing (a) and (b), our method also manages to typically/often stop sooner than Wald’s (though not guaranteeing (c)).
5.1 Two-sided stopping with valid Type-I error control
Suppose our test is based on a test (super)martingale $(M_{t})$, such as the likelihood ratio. We consider stopping for futility if $M_{t}\leq\nu_{t}$, where $\nu_{t}<1/\alpha$ is a predictable parameter, meaning $\nu_{t}$ is measurable with respect to $\mathcal{F}_{t-1}$. For example, Wald’s approximation suggests to set $\nu_{t}=\frac{\beta}{1-\alpha}$. Note that stopping for futility can never increase the type I error, but it may hurt the power. Below, we will show how we can gain back part of that power.
If $M_{t}>0$, we have $\mathbb{E}_{0}[M_{t^{\prime}}|\mathcal{F}_{t}]=M_{t}>0$ for all $t^{\prime}>t$. Hence, if we stop when our test martingale is strictly positive, we make a conservative decision since there is a positive probability of rejecting if we continue sampling. Therefore, our approach is to form a truncated test (super)martingale $(M_{t}^{\mathrm{trunc}})$ such that $M_{t}^{\mathrm{trunc}}\leq\nu_{t}$ implies that $M_{t}^{\mathrm{trunc}}=0$. To achieve this, we extend our truncation function $T_{\alpha}(x;M)$ from (8) as follows
$T_{\alpha}(x;M,\nu)=\begin{cases}0,&\text{ if }Mx\leq\nu\\ x,&\text{ if }\nu<Mx\leq\frac{1}{\alpha}\\ \frac{1}{M\alpha},&\text{ if }Mx>\frac{1}{\alpha}.\end{cases}$ (22)
Note that $T_{\alpha}(x;M,\nu)=T_{\alpha}(x;M)$, if $\nu=0$. Furthermore, $T_{\alpha}(x;M,\nu)\leq T_{\alpha}(x;M)$ for all $x\in\mathbb{R}_{\geq 0}$, implying that the boosting factors obtained by $T_{\alpha}(x;M,\nu)$ are larger than the boosting factors computed with $T_{\alpha}(x;M)$. We can use $T_{\alpha}(x;M,\nu)$ for boosting in the same manner as we did for $T_{\alpha}(x;M)$ in Algorithm 1. The following corollary extends Theorem 2.1 by including the futility stop.
Corollary 5.1.
Let $(M_{t})$ be any test (super)martingale and $(M_{t}^{\mathrm{boost}})$ be the process obtained by applying Algorithm 1 with $T_{\alpha}(x;M_{t-1}^{\mathrm{boost}},\nu_{t})$ and define the sequential tests $\delta_{M}=\mathbbm{1}\{M_{\tau_{M}}\geq 1/\alpha\}$ and $\delta_{\mathrm{boost}}=\mathbbm{1}\{M_{\tau_{\mathrm{boost}}}^{\mathrm{boost}}\geq 1/\alpha\}$, where $\tau_{M}=\inf\{t\in\mathbb{N}:M_{t}\geq 1/\alpha\lor M_{t}\leq\nu_{t}\}$ and $\tau_{\mathrm{boost}}=\inf\{t\in\mathbb{N}:M_{t}^{\mathrm{boost}}\geq 1/\alpha\lor M_{t}^{\mathrm{boost}}\leq\nu_{t}\}$. Then $M_{t}^{\mathrm{boost}}\geq M_{t}$ for all $t<\tau_{M}$ and $\delta_{\mathrm{boost}}\geq\delta$ almost surely. Furthermore, $(M_{t}^{\mathrm{boost}})$ is a test supermartingale. Therefore, $\mathbb{P}_{0}(\delta_{\mathrm{boost}}=1)\leq\alpha$ and $\mathbb{E}_{0}[M_{\tau}^{\mathrm{boost}}]\leq 1$ for all stopping times $\tau$.
In the same manner as in Equation (13), we can calculate for power-one SPRTs
$\mathbb{E}_{\theta_{0}}[T_{\alpha}(b_{t}\lambda(X_{t});\Lambda_{t-1}^{\mathrm{boost}},\nu_{t})|\Lambda_{t-1}^{\mathrm{boost}}]$ $=b_{t}\left[F_{\theta_{1}}\left(\lambda^{-1}\left([b_{t}\alpha\Lambda_{t-1}^{\mathrm{boost}}]^{-1}\right)\right)-F_{\theta_{1}}\left(\lambda^{-1}\left(\nu_{t}/[b_{t}\Lambda_{t-1}^{\mathrm{boost}}]\right)\right)\right]$
$+(\Lambda_{t-1}^{\mathrm{boost}}\alpha)^{-1}\left[1-F_{\theta_{0}}\left(\lambda^{-1}\left(\left[b_{t}\alpha\Lambda_{t-1}^{\mathrm{boost}}\right]^{-1}\right)\right)\right].$
The boosting factors obtained in the same setup as for Table 2, but with a stop for futility at $\nu_{t}=0.4$, is shown in Table 3. Obviously, all boosting factors are larger when the stop for futility is included. Furthermore, especially in cases where the current martingale value $\Lambda_{t-1}^{\mathrm{boost}}$ is small, including the stop for futility increases the boosting factor.
$\delta$ $\Lambda_{t-1}^{\mathrm{boost}}$ $0.5$ $1$ $2$ $4$ $10$
$0.1$ $1.00895$ $1$ $1$ $1$ $1.00001$
$0.5$ $1.17964$ $1.01743$ $1.00026$ $1.00019$ $1.03019$
$1$ $1.21801$ $1.07547$ $1.02817$ $1.05651$ $1.37357$
$2$ $1.32013$ $1.38991$ $1.62094$ $2.21769$ $5.76214$
$3$ $2.73073$ $3.75762$ $6.00201$ $12.1467$ $68.8295$
Table 3: Boosting factors obtained for $\alpha=0.05$ and $\nu_{t}=0.4$, $t\in\mathbb{N}$, in a simple Gaussian testing problem with mean difference $\delta$ and variance $1$.
Remark 3.
Note that our boosted SPRT could also be written as
$\text{reject }H_{0},\text{ if }\Lambda_{t}\geq\gamma_{1}^{t},\quad\text{accept}H_{0},\text{ if }\Lambda_{t}\leq\gamma_{0}^{t},\quad\text{and continue sampling if }\gamma_{0}^{t}<\Lambda_{t}<\gamma_{1}^{t},$
where $\gamma_{1}^{t}=\frac{1}{\alpha\prod_{i=1}^{t}b_{i}}$ and $\gamma_{0}^{t}=\frac{\nu_{t}}{\prod_{i=1}^{t}b_{i}}$ are predictable parameters. Such generalized SPRTs with non-constant boundaries over time were already introduced by Weiss [44], however, Weiss assumed that $\gamma_{0}^{t}$ and $\gamma_{1}^{t}$, $t\in\mathbb{N}$, are prespecified constants.
Remark 4.
Typically, the parameter $\nu_{t}$ is set to some value smaller than $1$, as $\Lambda_{t}<1$ implies that the current data favors the null hypothesis over the alternative. However, $\nu_{t}$ is a predictable parameter and one could also think of choosing it based on the current value of $\Lambda_{t-1}^{\mathrm{boost}}$. For example, it might be reasonable to set $\nu_{t}=8$ if $\Lambda_{t-1}^{\mathrm{boost}}=10$ to stop in case of decreasing evidence. In general, it is even possible to set $\nu_{t}>\Lambda_{t-1}^{\mathrm{boost}}$.
5.2 Two-sided stopping with valid type-I and type-II error control
We know that boosting always ensures type I error control at the level $\alpha$ by Ville’s inequality (5). Now we want to use the same approach to choose the stop for futility $\nu_{t}$ in a way that ensures valid type II error control at a level $\beta\in(0,1)$. The idea is to use a second SPRT for testing the alternative $H_{1}:X_{i}\sim\mathbb{P}_{1}$ against the null hypothesis $H_{0}:X_{i}\sim\mathbb{P}_{0}$ and stop for futility with our original test if this “inverse SPRT” rejects $H_{1}$. Hence, instead of using the likelihood ratio process $(\Lambda_{t})$ of $\mathbb{P}_{1}$ against $\mathbb{P}_{0}$, we use its inverse $\Lambda_{t}^{\mathrm{inv}}=1/\Lambda_{t}$. By Ville’s inequality (5), we control the type II error for the original test of $H_{0}$ against $H_{1}$ at level $\beta$, if we stop for futility when the inverse likelihood ratio $\Lambda_{t}^{\mathrm{inv}}$ crosses $1/\beta$. However, as we demonstrated in the previous sections, Ville’s inequality is loose if there is overshoot. Hence, to get tighter type II error control, we need to boost the inverse SPRT in the same manner as we did for the usual SPRT and include the stop for futility yielded by the inverse SPRT in the truncation function (22). This requires solving the following two-dimensional equation system at each step $t\in\mathbb{N}$ for the boosting factors $b_{t}$ and $b_{t}^{\mathrm{inv}}$,
[…]
with $\lambda_{t}(X_{t})=p_{1}(X_{t})/p_{0}(X_{t})$,
$\Lambda_{t}^{\mathrm{boost}}=\prod_{i=1}^{t}T_{\alpha}(b_{i}\lambda_{i}(X_{i});\Lambda_{i-1}^{\mathrm{boost}},\nu_{i})\quad\text{and}\quad\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}=\prod_{i=1}^{t}T_{\alpha}(b_{i}/\lambda_{i}(X_{i});\Lambda_{i-1}^{\mathrm{boost},\mathrm{inv}},\nu_{i}^{\mathrm{inv}}).$
The definition of $\nu_{t}$ ensures that the original SPRT only stops for futility when the boosted inverse SPRT stops for rejection. In the following proposition, we use this to prove that this indeed yields valid error control.
Proposition 5.2.
Let the boosting factors $b_{1},b_{2},\ldots$ and inverse boosting factors $b_{1}^{\mathrm{inv}},b_{2}^{\mathrm{inv}},\ldots$ be chosen such that (23) and (24) are fulfilled for each $t\in\mathbb{N}$. Then the boosted SPRT given by $\delta^{\mathrm{boost}}=\mathbbm{1}\{\Lambda_{\tau_{\mathrm{boost}}}^{\mathrm{boost}}\geq 1/\alpha\}$ with $\tau_{\mathrm{boost}}=\inf\{t\in\mathbb{N}:\Lambda_{t}^{\mathrm{boost}}\geq 1/\alpha\lor\Lambda_{t}^{\mathrm{boost}}\leq\nu_{t}\}$ controls the type I error at level $\alpha$ and the type II error at level $\beta$.
Proof.
Due to (23), the type I error control follows immediately by Ville’s inequality. Now, let $\tau_{\mathrm{inv}}=\inf\{t\in\mathbb{N}:\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\geq 1/\beta\lor\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\leq\nu_{t}^{\mathrm{inv}}\}$ be the stopping time of the inverse SPRT. To prove type II error control, we first show that $\tau_{\mathrm{boost}}=\tau_{\mathrm{inv}}$ almost surely. For this, note that
$\beta\prod_{i=1}^{t}b_{i}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}>1/\alpha\Leftrightarrow\alpha\prod_{i=1}^{t}b_{i}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}>1/\beta,$
and therefore $\nu_{t}<1/\alpha$ iff $\nu_{t}^{\mathrm{inv}}<1/\beta$. On the event $\nu_{1}<1/\alpha,\ldots,\nu_{t}<1/\alpha$, we have
$\Lambda_{t}^{\mathrm{boost}}\leq\nu_{t}\ \Leftrightarrow\ \Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\geq 1/\beta\quad\text{and}\quad\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\leq\nu_{t}^{\mathrm{inv}}\ \Leftrightarrow\ \Lambda_{t}^{\mathrm{boost}}\geq 1/\alpha.$
Furthermore, if $\nu_{t}=1/\alpha$ and we did not stop until step $t$ ($\tau_{\mathrm{boost}}\geq t$), we always have $\tau_{\mathrm{boost}}=t$ and $\tau_{\mathrm{inv}}=t$. Therefore, $\tau_{\mathrm{boost}}=\tau_{\mathrm{inv}}$. With this, it follows that
$\Lambda_{t}^{\mathrm{boost}}\leq\nu_{t}\land\tau_{\mathrm{boost}}=t$ $\implies\Lambda_{t}\leq\beta\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}\land\tau_{\mathrm{boost}}=t$
$\Leftrightarrow\Lambda_{t}^{\mathrm{inv}}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}\geq 1/\beta\land\tau_{\mathrm{boost}}=t$
$\Leftrightarrow\Lambda_{t}^{\mathrm{inv}}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}\geq 1/\beta\land\tau_{\mathrm{inv}}=t$
$\Leftrightarrow\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\geq 1/\beta\land\tau_{\mathrm{inv}}=t,$
where the last equivalence follows from the fact that if we have not stopped for futility yet, then $\Lambda_{t}^{\mathrm{inv}}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}\geq 1/\beta$ iff $\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\geq 1/\beta$. Due to Ville’s inequality and (24), the probability that there exists a $t\in\mathbb{N}$ satisfying the latter inequality is bounded by $\beta$ under $H_{1}$. ∎
Note that the error control of this strategy can still be slightly conservative, since it can happen that
$\beta\prod_{i=1}^{t}b_{i}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}>1/\alpha\Leftrightarrow\alpha\prod_{i=1}^{t}b_{i}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}>1/\beta,$
which means that the rejection region of the boosted SPRT and the boosted inverse SPRT overlap. For this reason, we set $\nu_{t}=1/\alpha$ and $\nu_{t}^{\mathrm{inv}}=1/\beta$ in this case which ensures valid type I and type II error control but avoids the simultaneous acceptance and rejection of both hypotheses. However, this conservatism is expected to be small, particularly if $\alpha$ and $\beta$ are small.
Figure 5: Comparison of sample size, type II and type I error probability between the SPRT with Wald’s approximated thresholds and the boosted SPRT in a simple Gaussian testing setup for different desired type II error probabilities ($\beta$) and $\alpha=0.01$. The boosted method provably controls the type I and type II error, while the SPRT with Wald’s approximated thresholds does not. Furthermore, the boosted SPRT saves sample size and nearly exhausts the error probabilities.
In Figure 5, we compare the sample size, the type II and the type I error probability of the SPRT with Wald’s approximated thresholds to our boosted method for different parameters $\beta$ in the same Gaussian simulation setup as considered in Section 3.1 (the type I error probability was obtained via importance sampling, see (15)). The desired type I error probability was set to $\alpha=0.05$ and the strength of the signal to $\delta=0.3$. The boosted method reduced the required sample size in all cases. Furthermore, the boosted SPRT nearly exhausts the type I error and type II error probability, particularly if $\beta$ is small, while the SPRT is conservative. Also note that the boosted SPRT provably controls the type I and type II error, but the SPRT with Wald’s approximated thresholds does not. In Appendix A we repeat this experiment for $\alpha=0.01$ instead of $\alpha=0.05$, which leads to similar results.
Remark 5.
In order to avoid solving the system of equations (23) and (24), one could also just set $\nu_{t}=\min\left(\beta\prod_{i=1}^{t}b_{i},1/\alpha\right)$ and solve for $b_{t}$ in
$\mathbb{E}_{0}\left[T_{\alpha}(b_{t}\lambda_{t}(X_{t});\Lambda_{t-1}^{\mathrm{boost}},\nu_{t})\Big{|}\mathcal{F}_{t-1}\right]=1.$
This would still lead to valid type I and type II error control, however, the type II error control would usually be more conservative than with the proposed approach.
Remark 6.
To ensure error control it is generally sufficient to satisfy (23) and (24) with inequalities $\leq$ (for all four equations). However, in order to obtain tight control the parameters $b_{t}$ and $b_{t}^{\mathrm{inv}}$ should be chosen as large as possible.
6 Discussion
6.1 Testing a forecaster
The classical SPRT by Wald [36, 37] assumes that the null hypothesis and the alternative hypothesis are specified in advance in the form of simple distributions $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$, respectively. However, in Section 3.2 we have seen that the SPRT remains even valid if we replace the fixed alternative $\mathbb{P}_{1}$ with a predictable plugin $\mathbb{P}_{1}^{t}$ that depends on the data $X_{1},\ldots,X_{t-1}\in\mathcal{X}$ observed so far. In general, the null distribution $\mathbb{P}_{0}^{t}$ and the alternative distribution $\mathbb{P}_{1}^{t}$ considered at each step $t\in\mathbb{N}$ can be chosen freely based on the previous data $X_{1},\ldots,X_{t-1}$ and some additional information $Z_{1},\ldots,Z_{t}\in\mathcal{Z}$. Shafer [24] formulated this as a forecaster-skeptic testing protocol, where the forecaster announces a probability distribution $\mathbb{P}_{0}^{t}$ at each step $t$ and a skeptic tests the forecaster by specifying an alternative $\mathbb{P}_{1}^{t}$. For example, one can think of the forecaster as a weather forecaster, which is tried to be invalidated by a skeptic. We summarized this general testing protocol in Algorithm 2.
The extension of our boosting technique (Algorithm 1) to the general testing problem in Algorithm 2 is straightforward, since at each step $t$ we only require the current martingale value $M_{t-1}=\Lambda_{t-1}$ and the distribution of the next factor $L_{t}=\frac{p_{1}^{t}(X_{t})}{p_{0}^{t}(X_{t})}$ under the null hypothesis $\mathbb{P}_{0}^{t}$, where $p_{0}^{t}$ and $p_{1}^{t}$ are the densities of $\mathbb{P}_{0}^{t}$ and $\mathbb{P}_{1}^{t}$, respectively.
Algorithm 2 SPRT in a general forecaster-skeptic testing protocol
Input: Lower and upper SPRT thresholds $\gamma_{0}$ and $\gamma_{1}$.
      Output: Decision on the forecaster.
$\Lambda_{0}=1$
2:for $t=1,2,\ldots$ do
3:     Reality announces $Z_{t}\in\mathcal{Z}$
4:     Forecaster announces a probability distribution $\mathbb{P}_{0}^{t}$ on $\mathcal{X}$
5:     Skeptic announces a probability distribution $\mathbb{P}_{1}^{t}$ on $\mathcal{X}$
6:     Reality announces $X_{t}\in\mathcal{X}$
7:     $\Lambda_{t}=\Lambda_{t-1}\cdot\frac{p_{1}^{t}(X_{t})}{p_{0}^{t}(X_{t})}$
8:     if $\Lambda_{t}\geq\gamma_{1}$ then
9:         Stop and reject the forecaster
10:     end if
11:     if $\Lambda_{t}\leq\gamma_{0}$ then
12:         Stop and accept the forecaster
13:     end if
14:end for
6.2 Other approaches to avoid Wald’s approximated thresholds
The problem of “overshoot” in the SPRT was already noted by Wald [36, 37] and examined in many previous works. One line of work approximates the exact type I and type II error of SPRTs based on Brownian motion and renewal theory [25, 27, 3, 13, 45, 17]. While the resulting approximations are asymptotically correct, they are very difficult to compute and therefore only used rarely in practice [25, 27]. For example, we have already seen that the available R packages for performing an SPRT [29, 2] rely on Wald’s approximated thresholds and that most applied studies also use them (e.g. [28, 14, 4, 10]). Another option is to truncate the SPRT by prespecifying a maximum number of observations. For simple distributions of the likelihood ratio and sufficiently small maximum numbers of observations it is then possible to calculate the type I and type II error for given thresholds exactly [1], which can be used to find appropriate bounds for the SPRT. Such an approach is frequently employed in group sequential trials [15, 41]. For more complex distributions one could approximate the error probabilities using Monte-Carlo sampling. This often leads to considerably computational effort, which sometimes can be reduced using importance sampling [26]. A limitation these three approaches have in common is that the used test statistics including their distributions must be specified in advance. This is not necessarily the case in the general forecaster-skeptic testing protocol described in the previous section. Hence, while our boosting approach is applicable to such complex and adaptive testing protocols, all of the aforementioned methods fail.
6.3 Improving (sequential) tests based on e-values
A recently popularized tool for sequential statistics is the e-value [20, 24, 9, 32]. An e-value for a null hypothesis $H_{0}:\mathbb{P}\in\mathcal{P}$, where $\mathcal{P}$ is some set of distributions, is a nonnegative random variable with expected value less or equal than one for all $\mathbb{P}\in\mathcal{P}$. For example, a likelihood ratio for a simple null hypothesis is an e-value. Therefore, e-values can be interpreted as generalized likelihood ratios for composite null hypotheses [24, 9]. When used for testing, e-values suffer from the same conservativeness due to overshoot as test martingales. It is conceptually straightforward to extend our boosting approach to general e-values theoretically by ensuring that (10) holds for all $\mathbb{P}\in\mathcal{P}$, although it might be difficult to calculate boosting factors in practice, depending on the complexity of $\mathcal{P}$. However, in Section 3.3 we at least showed that it is possible to apply it with the composite null hypotheses resulting from one-sided tests for exponential families.
7 Summary
In this paper, we introduced a general approach to uniformly improve nonnegative (super)martingale based sequential tests by avoiding an overshoot at the upper threshold $1/\alpha$ (where $\alpha$ is the desired type-I error). We illustrated its application by the power-one SPRT with one-parameter exponential families for one-sided null and alternative hypotheses. We also showed how it can be applied when a lower threshold is specified, that determines an early stopping time for accepting the null hypothesis. With this we obtain an alternative to the two-sided SPRT with Wald’s approximated thresholds, where our alternative provides valid type I and type II error control while often reducing the sample size. Our sequential tests can also be used to tighten confidence sequences and be extended to nontrivial settings like sampling without replacement and conformal martingales.
Acknowledgments
LF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project number 281474342/GRK2224/2. AR was funded by NSF grant DMS-2310718. We are grateful to Wouter M. Koolen for pointing out the possibility of using an inverted SPRT to control the type II error, and to Nick Koning for mentioning the log-optimality of boosting factors. In addition, we would like to thank Yajun Mei and Qunzhi Xu for a useful discussion.
Appendix A Additional simulation results
In this section, we provide some additional simulation result. In Figure 6 we repeated the simulated experiment for simple hypotheses from Figure 1 but with $\alpha=0.01$ instead of $\alpha=0.05$. Obviously, the required sample size is larger for $\alpha=0.01$ than for $\alpha=0.05$. It can also be seen that the gain in percentage of sample size is smaller than for $\alpha=0.05$, however, due to the larger total sample size the absolute gain in sample size remains similar.
Figure 6: Comparison of sample size and type I error between the SPRT and the boosted SPRT in a simple Gaussian testing setup for different signal strengths and $\alpha=0.05$. The boosted SPRT saves $2-8\%$ of sample size compared to the SPRT and exhausts the significance level. The SPRT is becoming increasingly conservative for stronger signals.
In Figure 7, we repeated the two-sided experiment from Figure 5 for $\alpha=0.01$. Again, the boosted SPRT saves a substantial number of samples compared to Wald’s two-sided SPRT with approximated thresholds and nearly exhausts the type I and type II error. Interestingly, the percentage of saved samples approximately increases with $\beta$ for $\alpha=0.01$, while it decreases for $\alpha=0.05$.
Figure 7: Comparison of sample size, type II and type I error probability between the SPRT with Wald’s approximated thresholds and the boosted SPRT in a simple Gaussian testing setup for different desired type II error probabilities ($\beta$) and $\alpha=0.01$. The boosted method provably controls the type I and type II error, while the SPRT with Wald’s approximated thresholds does not. Furthermore, the boosted SPRT saves sample size and nearly exhausts the error probabilities.